OpenStudy (nick1232):

-70.098=2.2(-7.3k-8)+0.5

OpenStudy (nick1232):

please help

OpenStudy (anonymous):

i am assuming you have to find the value of k, k = 3.3

OpenStudy (nick1232):

Can you show the work ?

OpenStudy (anonymous):

subtract 0.5 on both sides of the equal sign then divide both sides of the equal sign by 2.2 then add 8 to both sides of the equal sign and then divide both sides of the equal sign by -7.3

OpenStudy (nick1232):

Ok thanks

OpenStudy (anonymous):

-70.098=2.2(-7.3k-8)+0.5 -70.098 - 0.5=2.2(-7.3k-8) -70.598 = 2.2(-7.3k) - 2.2(8) -70.598 = -16.06k - 17.6 -52.998 = -16.06k 52.998 = 16.06k k = 3.3

OpenStudy (nick1232):

Thats better (:

OpenStudy (nick1232):

Thank you

OpenStudy (anonymous):

welcome:)